OpenStudy (anonymous):

Solve for the variable in this equation. 1/4(2x-12)=2/3(3x+9) x=

OpenStudy (anonymous):

\[1/4(2x-12)=2/3(3x+9)\] multiply by 12 to get \[3(2x-12)=8(3x+9)\] distribute to get \[6x-36=24x+72\] and it should be good from there

OpenStudy (anonymous):

subtract \(6x\) from both sides \[-36=18x+72\] subtract 72 from both sides \[-108=18x\]divide by 18 \[x=-6\]

OpenStudy (anonymous):

Oh so the answer is -6

OpenStudy (anonymous):

is it?